# Physics-aware Machine Learning Revolutionizes Scientific Paradigm for Machine Learning and Process-based Hydrology

### v4

@@ 4 Summary and Future Directions
PaML is a promising technique to take the best from both physics-based modeling and state-of-the-art ML models to improve the physical consistency, generalization, interpretability, and causality of ML. We present a comprehensive review of PaML methods, categorizing them into four aspects: physical data-guided ML, physics-informed ML, physics-embedded ML, and physics-aware hybrid learning. Advances in the PaML could promote ML-aided hypotheses in solving scientific problems, enabling not only rapidly exploiting the information in big data, but also scientific discoveries derived from new theorizations.
@@ Appendix A. Physical Methods in Hydrological Process
The hydrological model can be categorized into three main groups: empirical models, conceptual models, and physically-based models [2]. Empirical models utilize existing data without accounting for the specific features and processes of hydrological systems, and are therefore referred to as data-driven models. On the other hand, physical and conceptual models require a comprehensive understanding of the physics governing the movement of surface water in the hydrological cycle [316].
@@
Conceptual models describe runoff processes by linking simplified components of the hydrological cycle using interconnected storages to represent different components. These models are typically lumped and employ uniform parameter values across the entire watershed. They necessitate a variety of parameters and meteorological input data. Calibration of these models requires a substantial amount of meteorological and hydrological records. A host of conceptual models have been proposed in the past including HBV model [317], non-recorded catchment areas (NRECA) [318], Boughton model [319], ARNO model [320], Xinanjiang model [321], etc.
@@
Physical-based models are grounded in an understanding of the physics governing hydrological processes. These models utilize fundamental physical laws and principles, including equations for water balance, conservation of mass and energy, momentum, and kinematics. Typically, finite difference or finite computation schemes are employed to solve these equations. For being built on physical laws and equations, spatially-explicit, process-based models are, in theory, capable to respond correctly to dynamic changes in catchment conditions and climate forcing; they can thus be used for the assessment of non-linearities between states, drivers, and catchment responses [322, 323]. Therefore, they are practical tools to predict, project, and attribute impacts of global change (e.g. climate change, land use change, water management) on the water cycle and even hydrological extremes. The primary advantage of physical models lies in their ability to connect model parameters with physical characteristics of the catchment, thereby enhancing their realism. Since physical models are semi- or fully distributed across watersheds, they necessitate a significant amount of data such as soil-hydraulic properties, vegetation, land cover characteristics, initial water depth, topography, and dimensions of the river network. This data is crucial for investigating changes in the hydrological cycle resulting from human activities and climate change, which are vital for effective water resources management [324]. Some examples of physical models include soil and water assessment tool (SWAT) [325], system hydrologique european (SHE) [326] or MIKE SHE [327], visualizing ecosystem land management assessments (VELMA) [328], kinematic runoff and erosion model (KINEROS) [321], Penn State integrated hydrologic modeling system (PIHM) [329], variable infiltration capacity (VIC) [330], and others.
@@ Appendix B. Physical Methods in Hydrodynamic Modeling
1D models. 1D hydrodynamic models have been widely used in modeling water flows [334]. Many hydraulic situations can make the 1D assumption, either because a more detailed solution is unnecessary (e.g., the application does not necessitate knowledge in other dimensions) or because the flow is distinctly 1D, such as in a confined channel or in a channelized flow. Typically, 1D models solve equations derived to ensure conservation of mass and momentum, resulting in the well-known 1D Saint-Venant equations,
@@
2D models. The 2D models depict water or flood flow as a two-dimensional field with the assumption that the third dimension, water depth, is shallow compared to the other two dimensions [336, 337]. Most methods solve two-dimensional shallow water equations, which describe conservation of mass and momentum in a plane and are derived by depth-averaging the Navier-Stokes equations,
@@
$\begin{gathered}\frac{\partial h}{\partial t}+\frac{\partial q_{x}}{\partial x}+\frac{\partial q_{y}}{\partial y}=0\\ \underbrace{\frac{\partial q_{x}}{\partial t}}_{\begin{subarray}{c}\text{ local }\\ \text{ acceleration }\end{subarray}}+\underbrace{\frac{\partial}{\partial x}\left(uq_{x}\right)+\frac{\partial}{\partial y}\left(vq_{x}\right)}_{\begin{array}[]{c}\text{ convective }\\ \text{ acceleration }\end{array}}+\underbrace{gh\frac{\partial(h+z)}{\partial x}}_{\begin{array}[]{c}\text{ pressure }+\\ \text{ bed gradients }\end{array}}+\underbrace{\frac{gh^{2}\|\mathbf{q}\|q_{x}}{h^{7/3}}}_{\text{friction }}=0\\ \underbrace{\frac{\partial q_{y}}{\partial t}}_{\begin{array}[]{c}\text{ local}\\ \text{ acceleration }\end{array}}+\underbrace{\frac{\partial}{\partial y}\left(vq_{y}\right)+\frac{\partial}{\partial x}\left(uq_{y}\right)}_{\begin{array}[]{c}\text{ convective }\\ \text{ acceleration }\end{array}}+\underbrace{gh\frac{\partial(h+z)}{\partial y}}_{\begin{array}[]{c}\text{ pressure }+\\ \text{ bed gradients }\end{array}}+\underbrace{\frac{gn^{2}\|\mathbf{q}\|q_{y}}{h^{7/3}}}_{\text{friction }}=0,\end{gathered}$ (4)
@@
here, $x$ and $y$ represent the two Cartesian directions; $q_{x}$ and $q_{y}$ denote the $x$ and $y$ components of the discharge per unit width vector $\mathbf{q}$; $u$ and $v$ are the $x$ and $y$ components of the flow velocity, respectively; $z$ denotes the bed elevation; and $n$ represents the Manning’s friction coefficient. In flood situations, the significant simplification introduced by the local inertial approximation relies on the assumption that the convective acceleration terms are negligible compared to other terms and can thus be disregarded.
@@
3D models. For various water flow scenarios, there has been a tendency to consider the detailed representation of flow dynamics in 3D unnecessary, favoring the adequacy of a 2D shallow water approximation [338]. Nevertheless, it is crucial to model vertical turbulence, vortices, and floods caused by dam breaks. Consequently, 3D models have been developed to incorporate vertical features. Some of these models solve horizontal flow using 2D shallow water equations and incorporate a quasi-three-dimensional extension to represent velocity in vertical layers. Other 3D models are based on the 3D Navier-Stokes equations, which describe the movement of fluid substances and are typically formulated as,
$\begin{gathered}\frac{\partial u}{\partial t}+u\cdot\nabla u+\frac{1}{\rho}\nabla p=g+\mu\nabla\cdot\nabla u,\\ \quad\nabla\cdot u=0,\end{gathered}$ (5)
@@
where $u$ represents velocity; $\rho$ denotes fluid density; $p$ stands for pressure; $g$ is gravitational acceleration; and $\mu$ represents kinematic viscosity. The incompressibility condition ($\quad\nabla\cdot u=0$) assumes that material density remains constant within a fluid parcel. Depending on how the processes are represented, these models can be categorized into two main types: Eulerian (grid-based) and Lagrangian (particle-based).
Acknowledgements
The work is jointly supported by the Helmholtz Association through the Framework of the Helmholtz Excellent Professorship “Data Science in Earth Observation - Big Data Fusion for Urban Research”(grant number: W2-W3-100), by the German Federal Ministry of Education and Research (BMBF) in the framework of the international future AI lab "AI4EO – Artificial Intelligence for Earth Observation: Reasoning, Uncertainties, Ethics and Beyond" (grant number: 01DD20001) and by German Federal Ministry for Economic Affairs and Climate Action in the framework of the "national center of excellence ML4Earth" (grant number: 50EE2201C).

### v5

@@ 3.3 Proposed HydroPML-based Application Highlights
Although these application cases primarily focus on the physical processes of water bodies, HydroPML can also be extended to other areas of hydrology, such as water quality. Recently, DL has become increasingly reliable for understanding and predicting water quality [328, 329], such as water temperature and dissolved oxygen dynamics [330, 331]. However, most approaches focus on water quality management and prediction under the PDgML frameworks. By integrating the physics of water-quality dynamics into ML frameworks (PeML), optimization processes (PiML), and hybrid learning (PaHL), the data requirements for ML can be reduced. This provides the possibility of achieving effective water-quality predictions for extreme events. Additionally, physic-discovery NNs can facilitate the discovery of new knowledge in water-quality dynamics [329]. Ultimately, HydroPML offers a pathway for integrating data-driven ML methods and physics-based models in hydrology.
@@ 4 Summary and Future Directions
PaML is a promising technique to take the best from both physics-based modeling and state-of-the-art ML models to improve the physical consistency, generalization, interpretability, and causality of ML. We present a comprehensive review of PaML methods, categorizing them into four aspects: physical data-guided ML, physics-informed ML, physics-embedded ML, and physics-aware hybrid learning. Advances in the PaML could promote ML-aided hypotheses in solving scientific problems, enabling not only rapidly exploiting the information in big data. but also scientific discoveries derived from new theorizations.
@@
Acknowledgements
The work is jointly supported by the German Federal Ministry of Education and Research (BMBF) in the framework of the international future AI lab "AI4EO – Artificial Intelligence for Earth Observation: Reasoning, Uncertainties, Ethics and Beyond" (grant number: 01DD20001) and the project Inno__\__MAUS (grant number: 02WEE1632B), by German Federal Ministry for Economic Affairs and Climate Action in the framework of the "national center of excellence ML4Earth" (grant number: 50EE2201C), by the Excellence Strategy of the Federal Government and the Länder through the TUM Innovation Network EarthCare and by Munich Center for Machine Learning.
@@ Acknowledgements
Supplementary Table 1: Proposed Physical Data-guided Machine Learning.
Categories Types Examples Objective(s) Limitation(s)
Informer [49] Long sequence time-series forecasting
ConvLSTM [53] A spatio-temporal sequence forecasting problem
DisasterNets [54] Embedding machine learning in disaster mapping
Geometric deep learning [62] A spatially-irregular forecast method
Neural message passing-based models [71, 73, 72, 75, 76, 77] A general framework for supervised learning on graphs (arbitrary domains) Mesh-based or particle-based; message-passing between small-scale moving and interacting objects
Fourier neural operator (FNOs) [82] Geo-FNO [90], Domain agnostic FNO [91] Solve PDEs on arbitrary geometries Unstable, limited generalization, the basis functions in the Fourier space are generally frequency localised with no spatial resolution
Physics-discovery Neural Networks Physics-discovery with unavailable prior knowledge Symbolic neural networks [97, 98, 332] Infer concise equations without any prior knowledge Challenging task due to complex and nonlinear systems, as well as noisy and incomplete data
AI Feynman [99, 100] Discover generalized symmetries in complex data
GNN [101] Identify the nontrivial relation due to its excellent inductive capability
PDE-FIND [103] Data-driven discovery for PDEs
Supplementary Table 2: Proposed Physics-informed Machine Learning.
Categories Types Examples Objective(s) Limitation(s)
CINN [115] A simple and efficient characteristic-informed neural networks for solving forward and inverse problems in hyperbolic PDEs Limited application; limited expression and generalization ability
PI-PINN [125] Orthogonal polynomials are used to construct the NN for solving nonlinear partial differential equations Limited expressivity of framework
Improving speed and efficiency Meta-learning based PINNs de Avila Belbute-Peres et al. [130] meta-trains a hyper-network that for each task can generate weights for a small NN. Psaros et al. [131] meta-learns a loss function that is used to optimize the NN. Penwarden et al. [132] proposes a meta-learning approach that learns an initialization of weights such that the NN can be optimized quickly. Limited expressivity of framework, unsteady in solving related PDEs
PINN-SR [146] Physics-informed neural network with sparse regression to discover governing partial differential equations from scarce and noisy data for nonlinear spatio-temporal systems High computational cost; requiring a vast number of collocation points for complex systems; limitations of PINNs
Physics- informed Neural Operators WNO-based Physics-informed WNO [152] A physics-informed WNO for learning solution operators of families of parametric PDEs without labeled training data, leveraging the time-frequency localization advantage of wavelets Complexity; the absence of a standardized procedure to obtain the optimum and desired network architecture of a physics-informed WNO and to assign the weights of boundary/initial condition loss; the lack of information regarding how to learn the operator for systems involving multi-scale physics systems
Supplementary Table 3: Proposed Physical Equation-embedded Machine Learning.
MeshGraphNets [156] Mesh-based partial differential equations A framework for learning mesh-based simulations using graph neural networks: the mesh-space, spanned by the simulation mesh, and the Euclidean world-space in which the simulation manifold is embedded High computational complexity; hyperparameters
FGN [157] A Lagrangian viewpoint of the Navier-Stokes equation Fluid graph networks (FGN) uses graphs to represent the fluid field. In FGN, fluid particles are represented as nodes and their interactions are represented as edges. FGN decomposes the simulation scheme into separate parts, including advection, collision, and pressure projection Limited generalization and computational efficiency
Differentiable modules Differential equations These modules implement physical equations as differentiable feature space or computational graphs, enabling the optimization of dynamic processes with analytical gradients and therefore improving sample efficiency. Examples include differentiable physics-informed graph networks [161], a differentiable layer (PDE-Constrained-Layer) [162], physics-aware difference graph networks [163], physics-based finite difference convolution connection [164], and differentiable modeling [13] Limitations on large-scale applications; require an in-depth understanding of differential equations
@@ Appendix A. Physical Methods in Hydrological Process
The hydrological model can be categorized into three main groups: empirical models, conceptual models, and physically-based models [2]. Empirical models utilize existing datasets without accounting for the specific features and processes of hydrological systems, and are therefore referred to as data-driven models. On the other hand, physical and conceptual models require a comprehensive understanding of the movement of surface water in the hydrological cycle [333].
@@
Conceptual models describe runoff processes by linking simplified components of the hydrological cycle using interconnected storages to represent different components. These models are typically lumped and employ uniform parameter values across the entire watershed. They need a variety of parameters and meteorological input data. Calibration of these models requires a substantial amount of meteorological and hydrological records. A host of conceptual models have been proposed in the past including HBV model [334], non-recorded catchment areas (NRECA) [335], Boughton model [336], ARNO model [337], Xinanjiang model [338], etc.
@@
Physical-based models are grounded in an understanding of the physics governing hydrological processes. These models utilize fundamental physical laws and principles, including equations for water balance, conservation of mass and energy, momentum, and kinematics. Typically, finite difference or finite computation schemes are employed to solve these equations. For being built on physical laws and equations, spatially-explicit, process-based models are, in theory, capable to respond correctly to dynamic changes in catchment conditions and climate forcing; they can thus be used for the assessment of non-linearities between states, drivers, and catchment responses [339, 340]. Therefore, they are practical tools to predict, project, and attribute impacts of global change (e.g. climate change, land use change, water management) on the water cycle and even hydrological extremes. The primary advantage of physical models lies in their ability to connect model parameters with physical characteristics of the catchment, thereby enhancing their realism. Since physical models are semi- or fully distributed across watersheds, they need a significant amount of data such as soil-hydraulic properties, vegetation, land cover characteristics, initial water depth, topography, and dimensions of the river network. This data is crucial for investigating changes in the hydrological cycle resulting from human activities and climate change, which are vital for effective water resource management [341]. Some examples of physical models include soil and water assessment tool (SWAT) [342], system hydrologique european (SHE) [343] or MIKE SHE [344], visualizing ecosystem land management assessments (VELMA) [345], kinematic runoff and erosion model (KINEROS) [338], Penn State integrated hydrologic modeling system (PIHM) [346], variable infiltration capacity (VIC) [347], and others.
@@ Appendix B. Physical Methods in Hydrodynamic Modeling
1D models. 1D hydrodynamic models have been widely used in modeling water flows [351]. Many hydraulic situations can make the 1D assumption, either because a more detailed solution is unnecessary (e.g., knowledge in other dimensions is not required for the purpose) or because the flow is distinctly 1D, such as in a channelized flow or a confined channel. Typically, 1D models solve equations derived to ensure conservation of mass and momentum, resulting in the well-known 1D Saint-Venant equations,
@@
2D models. The 2D models depict water or flood flow as a two-dimensional field with the assumption that the third dimension, water depth, is shallow compared to the other two dimensions [353, 354, 35]. Most methods solve two-dimensional shallow water equations, which describe conservation of mass and momentum in a plane and are derived by depth-averaging the Navier-Stokes equations,
@@
$\begin{gathered}\frac{\partial h}{\partial t}+\frac{\partial q_{x}}{\partial x}+\frac{\partial q_{y}}{\partial y}=0,\\ \underbrace{\frac{\partial q_{x}}{\partial t}}_{\begin{subarray}{c}\text{ local }\\ \text{ acceleration }\end{subarray}}+\underbrace{\frac{\partial}{\partial x}\left(uq_{x}\right)+\frac{\partial}{\partial y}\left(vq_{x}\right)}_{\begin{array}[]{c}\text{ convective }\\ \text{ acceleration }\end{array}}+\underbrace{gh\frac{\partial(h+z)}{\partial x}}_{\begin{array}[]{c}\text{ pressure }+\\ \text{ bed gradients }\end{array}}+\underbrace{\frac{gh^{2}\|\mathbf{q}\|q_{x}}{h^{7/3}}}_{\text{friction }}=0,\\ \underbrace{\frac{\partial q_{y}}{\partial t}}_{\begin{array}[]{c}\text{ local}\\ \text{ acceleration }\end{array}}+\underbrace{\frac{\partial}{\partial y}\left(vq_{y}\right)+\frac{\partial}{\partial x}\left(uq_{y}\right)}_{\begin{array}[]{c}\text{ convective }\\ \text{ acceleration }\end{array}}+\underbrace{gh\frac{\partial(h+z)}{\partial y}}_{\begin{array}[]{c}\text{ pressure }+\\ \text{ bed gradients }\end{array}}+\underbrace{\frac{gn^{2}\|\mathbf{q}\|q_{y}}{h^{7/3}}}_{\text{friction }}=0,\end{gathered}$ (4)
@@
where the two Cartesian directions are represented by $x$ and $y$; $q_{x}$ and $q_{y}$ denote the $x$ and $y$ components of the discharge per unit width vector $\mathbf{q}$; $u$ and $v$ are the $x$ and $y$ components of the flow velocity, respectively; $z$ denotes the bed elevation; and $n$ represents the Manning’s friction coefficient. In flood situations, the significant simplification introduced by the local inertial approximation is based on the assumption that the convective acceleration terms are negligible compared to other terms, allowing them to be disregarded.
@@
3D models. For various water flow scenarios, there has been a tendency to consider the detailed representation of flow dynamics in 3D [355]. For example, it is crucial to model vertical turbulence, vortices, and floods caused by dam breaks. Consequently, 3D models have been developed to incorporate vertical features. Some of these models solve horizontal flow using 2D shallow water equations and incorporate a quasi-three-dimensional extension to represent velocity in vertical layers. Other 3D models are based on the 3D Navier-Stokes equations, which describe the movement of fluid and are typically formulated as,
$\begin{gathered}\frac{\partial u}{\partial t}+u\cdot\nabla u+\frac{1}{\rho}\nabla p=g+\mu\nabla\cdot\nabla u,\\ \nabla\cdot u=0,\end{gathered}$ (5)
@@
where $u$ represents velocity; $\rho$ denotes fluid density; $p$ stands for pressure; $g$ is gravitational acceleration; and $\mu$ represents kinematic viscosity. The incompressibility condition ($\nabla\cdot u=0$) assumes that material density remains constant within a fluid parcel. Based on the process representation, these models can be categorized into two primary types: grid-based Eulerian models and particle-based Lagrangian models.